OpenStudy (lgbasallote):

Note: This is NOT a question. This is a tutorial. How to take the LCD/LCM of a rational expression/number? See comment below to see how!

jhonyy9 (jhonyy9):

???

OpenStudy (lgbasallote):

Taking the LCD/LCM only involves a few steps. For the simplicity of this tutorial, I'll use fractions. Suppose you are given: \(\Large \frac{2}{3} + \frac{3}{4} + \frac{5}{6} + \frac{1}{9} + \frac{3}{8}\) STEP 1: Express all your denominators as factors of prime numbers/expressions. \(\Large \frac{2}{3} + \frac{3}{(2)(2)} + \frac{5}{(2)(3)} + \frac{1}{(3)(3)} + \frac{3}{(2)(2)(2)}\) Step 2: List down your factored denominators. Denominator 1: 3 Denominator 2: (2)(2) Denominator 3: (2)(3) Denominator 4: (3)(3) Denominator 5: (2)(2)(2) Step 3: Pick a number from this list. This becomes ONE of the factors of my LCD/LCM. Let's say I choose 3. My LCD/LCM now is \(3 \times a \times b \times c\) and so on, and so forth. Step 4: Cancel ONE of that number (the number you chose in Step 3, which in our case it's 3) from EACH your denominators, if possible. And so, we'll cancel 3 from each of our denominators. Denominator 1: \(\cancel{3}\) Denominator 2:(2)(2) <---- no 3 so we don't cancel anything Denominator 3: (2)\(\cancel{(3)}\) Denominator 4: (3)\(\cancel{(3)}\) <-----note that we only cancel one 3 Denominator 5: (2)(2)(2) <----again no 3 so we don't cancel anything. We repeat Step 3 and choose another number..this time I choose 2. So my LCD is now 3 x 2 but that's not yet finished. Again, we list down the remaining denominators. Denominator 1: 1 Denominator 2: \(\cancel{(2)}\)(2) <----note that we only cancel one 2 Denominator 3: \(\cancel{(2)}\)(1) Denominator 4: 3 Denominator 5: \(\cancel{(2)}\)(2)(2) We again repeat Step 3. This time, I'll choose 3 again. So our LCD so far is 3 x 3 x 2. Now, we list te remaining denominators. Denominator 1: 1 Denominator 2: 2 Denominator 3: 1 Denominator 4: \(\cancel{(3)}\) Denominator 5: (2)(2) Repeat Step 3 again. I'll choose 2 again this time. So our LCD is 3 x 3 x 2 x 2. List down the remaining denominators... Denominator 1: 1 Denominator 2: \(\cancel{(2)}\) Denominator 3: 1 Denominator 4: 1 Denominator 5: \(\cancel{(2)}\)(2) Repeat Step 3. The only remaining number is 2 so that is the last input to out LCD. So the LCD of \(\Large \frac{2}{3} + \frac{3}{4} + \frac{5}{6} + \frac{1}{9} + \frac{3}{8}\) is 3 x 3 x 2 x 2 x 2 = 9 x 8 = 72 <---LCD

OpenStudy (lgbasallote):

lol @jhonyy9 i had to type it :p =)))

jhonyy9 (jhonyy9):

yes is nice but very long - is more easy if you check step by step like - 1, for 3 and 4 you get 12 - so 12 is right for 6 too - for 12 and 9 make 12=3*4 and 9=3*3 so than will be right 3*3*4 = 36 - than for 36 and 8 so 36=3*3*2*2 and 8=2*2*2 so than we get 3*3*2*2*2 = =9*8=72

OpenStudy (lgbasallote):

hmm i see...i admit that is faster...though i just want my tutorials to be beginner friendly haha you know...see it immediately...though that is a cool method too :D

OpenStudy (thomas5267):

A great tutorial! Perhaps using the calculator to do prime decomposition will be quicker.

OpenStudy (lgbasallote):

haha thanks :) prime decomposition is fast...but i dont think it works on rational expressions? this does...i just used fractions for the simplicity of the method and to conserve latex haha

OpenStudy (lgbasallote):

jhony's method too..i dont think works on rational expressions...does it @jhonyy9 ?

jhonyy9 (jhonyy9):

can you writing an example for what not is right ?

OpenStudy (lgbasallote):

uhmm i dont know..i was just asking if your method works on rational expressions as well?

OpenStudy (zepp):

Haha, my method is almost exactly the same thing, here: http://puu.sh/slTq

OpenStudy (lgbasallote):

maybe i'll write an example of a rational expression... \(\Large \frac{4}{3x} + \frac{2x +5}{9x^2 + 18x + 9} + \frac{1}{9}\) using my method..we list down the denominators in factored form... \(\Large \frac{4}{(3)(x)} + \frac{2x +5}{(3)(3)(x+ 1)^2} + \frac{1}{(3)(3)}\) List the denominators... 1st denom -> (3)(x) 2nd denom - > (3)(3)(x+1)^2 3rd denom -> (3)(3) i'll choose 3.. cancel one 3 from each denom... 1st denom -> x 2nd denom -> 3(x+1)^2 3rd denom -> 3 i'll choose another 3..so my LCD so far is 9...cancel another 3 from the denoms... 1st denom -> x 2nd denom -> (x+1)^2 3rd denom -> 1 now i'll choose x (so LCD is now 9x)..cancel one x from each denom 1st denom -> 1 2nd denom -> (x+1)^2 3rd denom -> 1 sin (x+1)^2 is the only one left it automatically jons the LCD...and so..the combined LCD is 9x(x+1)^2

OpenStudy (lgbasallote):

i cut it short of explanation hope it's still understandable

jhonyy9 (jhonyy9):

yes right too

OpenStudy (zepp):

http://puu.sh/slVK :D

OpenStudy (lgbasallote):

oh hmm guess that works there too

OpenStudy (zepp):

Good job! ;D

OpenStudy (dumbcow):

very nice..they should have a seperate section for tutorials, that way your work is more permanent...since many users may not see this post

OpenStudy (dumbcow):

or write up your tutorial somewhere and then you could just post the link to OS periodically to save you time

OpenStudy (lgbasallote):

yeah i sometimes reference myself when answering questions :D as for the section...i just think a special status would be nice..because creating another section..we cant guarantee many people would see it either...maybe the admins can notice my works someday then they can think of something :D i think these have been brought up to the admin's attention already

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

Nice new profile pic lgb, I always knew you were a girl!

OpenStudy (lgbasallote):

kind of irrelevant to the thread but thanks :)

OpenStudy (callisto):

I agree with @dumbcow . Solving problems/ Answering questions is not like tutorial. Tutorials are more permanent. Sorry, that's irrelevant to the thread again :|

OpenStudy (lgbasallote):

haha no...that's relevant bc it's about the tutorial..not abt the topic but what the heck haha...well i guess we can wait for the admin's decision on this..i mean if im the only one doing this i dont think ill get granted that request :p ahaha i dunno

hero (hero):

This belongs in the Education Section.

OpenStudy (lgbasallote):

we have that? o.O well it's still math :p i think this is what they call "gray area"

hero (hero):

It should be created pretty soon actually. And you could declare your support for it in the next Open House

OpenStudy (lgbasallote):

hmm sure :DDD the open house isnt going to be a thread in the os feedback again is it?

hero (hero):

I don't know what they're going to do this time. @cshalvey would know better than I.

OpenStudy (anonymous):

lgbasa is having graet tutorial classes nowa days..

OpenStudy (lgbasallote):

haha thanks for that @shruti :) i try my best ^_^